are these equations parallel
x+11=y and x-y=-7

Answer by richwmiller(17219)

No. Equations are not parallel.
The lines represented by the equations may be parallel.
let's put them in slope intercept form
y=x+11
has a slope of 1
x-y=-7
add 7 and y to both sides
y=x+7
slope is also 1
So the lines are a parallel.